improper integrals
upper is infinity
lower 4
x^(-3/2) dx
having trouble can someone help please

Answer by Alan3354(21611)

upper is infinity
lower 4
x^(-3/2) dx
-----------
Int = x^(-1/2)/(-1/2)
= -2x^(-1/2) + C
At inf = 0 + C
At 4 = -1 + C
-------------
Diff = 1